The capacities (in ampere-hours) were measured for a sample of 120 batteries. The average was 178 and the standard deviation was 15.

a)Find a 95% confidence interval for the mean capacity of batteries produced by this method. Round the answers to three decimal places.The 95% confidence interval is?

b) Find a 99% confidence interval for the mean capacity of batteries produced by this method. Round the answers to three decimal places.The 99% confidence interval is?

c) An engineer claims that the mean capacity is between 175.864 and 180.136 ampere-hours. With what level of confidence can this statement be made? Express the answer as a percent and round to two decimal places.

d) Approximately how many batteries must be sampled so that a 95% confidence interval will specify the mean to within ±2 ampere-hours? Round up the answer to the nearest integer.

e) Approximately how many batteries must be sampled so that a 99% confidence interval will specify the mean to within ±2 ampere-hours? Round up the answer to the nearest integer

please dont answer the quastion if you dont know the right answer.
Answer #1

a)

sample mean sample size population std deviation X 178.000 120 15 nE standard errror of means ?,-o/vn 1.369 for 95 % CI value of margin of error E-z*std error lower confidence bound-sample mean-margin of error- Upper confidence bound-sample mean +margin of error 1.960 2.684 175.316 180.684

95% confidence interval is 175.316 ; 180.684

b)

for 99 % CI value of z= 2.580
margin of error E=z*std error                            = 3.533
lower confidence bound=sample mean-margin of error= 174.467
Upper confidence bound=sample mean +margin of error= 181.533

99% confidence interval is 174.467 ; 181.533

c)

here margin of error =(upper limit -lower limit)/2=(180.136-175.864)/2=2.136

therefore critical value z =margin of error/std error=2.136/1.369=1.56

from table 1.56 is critical value for 88% confidence interval

therefore level of confidence =88.00%

d)

for 95 % CI value of z= 1.960
standard deviation \sigma = 15.00
margin of error E = 2
required sample size n=(z\sigma/E)2    = 217.0

e)

for 99 % CI value of z= 2.580
required sample size n=(z\sigma/E)2    = 375